\begin{thebibliography}{99}
\small
  \bibitem{BM02} Bartlett, P.L., Mendelson, S.: Rademacher and gaussian complexities: risk bounds and structural results. J. Mach. Learn. Res. {\bf3}, 463-482 (2002)
       \bibitem{BJ11}  Brooks, J.P.: Support vector machines with the ramp loss and the hard margin loss. Oper. Res. {\bf59}(2), 467-479 (2011)
 \bibitem{CN14} Carrizosa, E., Nogales-Gmez, A., Romero, M.D.: Heuristic approaches for support vector machines with the ramp loss. Optim. Lett. {\bf8}(3), 1125-1135 (2014)



%   \bibitem{CS04} Christmann, A., Steinwart, I.: On robustness properties of convex risk minimization methods for pattern recognition.  J. Mach. Learn. Res. {\bf5}, 1007-1034 (2004)
           \bibitem{RF06a} Collobert, R., Sinz, F., Weston, J., Bottou, L.: Trading convexity for scalability. In: Proceedings of the 23rd International Conference on Machine Learning, pp. 201-208 (2006)

                    \bibitem{RF06b}  Collobert, R., Sinz, F., Weston, J., Bottou, L.: Large scale transductive SVMs. J. Mach. Learn. Res. {\bf7}, 1687-1712 (2006)

 \bibitem{CV95} Cortes, C., Vapnik, V.: Support vector networks. Mach. Learn. {\bf20}(3), 273-297 (1995)
%                                        \bibitem{CS13} Cotter, A., Shalev-Shwartz, S., Srebro, N: Learning optimally sparse support vector machines. In: International Conference on Machine Learning, pp. 266-274  (2013)
       \bibitem{EB11} Ertekin, S., Bottou, L., Giles, C.L.: Nonconvex online support vector machines. 	IEEE Trans. Pattern Anal. Mach. Intell. {\bf33}(2), 368-381 (2011)
            \bibitem{GG19}Gaudioso, M., Gorgone, E., Hiriart-Urruty, J.B.: Feature selection in SVM via polyhedral k-norm. Optim. Lett. {\bf14}(1), 19-36 (2020)
 \bibitem{HS14}Huang, X.L., Shi, L., Suykens, J.A.K.: Ramp loss linear programming support vector machine. J. Mach. Learn. Res. {\bf15}, 2185-2211 (2014)
%     \bibitem{HSS2014} Huang X L, Shi L, Suykens J A K. Support vector machine classifier with pinball loss. \emph{IEEE Transactions on Pattern Analysis and Machine Intelligence}, 36(5):984-997. (2014)

% \bibitem{GM14} Ghosh S, Mondal S, Ghosh B. A comparative study of breast cancer detection based on SVM and MLP BPN classifier.\emph{
%International Conference on First International Conference on Automation, Control, Energy
%and Systems}, 87-90. (2014) %(2014)}
% \bibitem{KK12}Khan N M, Ksantini R, Ahmad I S, et al. A novel SVM+NDA model for classification with an application to face recognition. \emph{Pattern Recognition}, 45(1):66-79. (2012)
% \bibitem{LC11}Li, C.H., Kuo, B.C., Lin, C.T., Huang, C.S.: A spatial-contextual support vector machine for remotely sensed image classification. IEEE Transactions on Geoence and Remote Sensing, {\bf50}(3), 784-799 (2012)
  \bibitem{KS09}Khemchandani, R., Suresh, C.: Optimal kernel selection in twin support vector machines. Optim. Lett. {\bf3}(1), 77-88 (2009)
      \bibitem{LS05}Liu, S.J., Shen, X.T., Wong, W.H.: Computational developments of $\psi$-learning. In: Proceedings of the 2005 SIAM International Conference on Data Mining, pp. 1-11 (2005)
%    \bibitem{LR20} Liu, Y.Y., Ren, L., Cao, X.H., Tong, Y.: Breast tumors recognition based on edge feature extraction using support vector machine.
%   Biomed. Signal Process. Control. https://doi.org/10.1016/j.bspc.2019.101825. (2020)
\bibitem{LZ20}  Liu, Z.Y., Kan, H.P., Zhang, T., Zhou, Z.J.: DUKMSVM: A framework of deep uniform kernel mapping support vector machine for short text classification.  Applied Sciences. {\bf10}(7), 1-16 (2020)

        \bibitem{LW19} Lu, S.J., Wang, H.R., Zhou, Z.J.: All-in-one multicategory ramp loss maximum margin of twin spheres support vector machine. Appl. Intell., {\bf49}(6), 2301-2314 (2019)
    \bibitem{DE15}  Mancev, D.: A sequential dual method for the structured ramp loss minimization. Facta Universitatis, {\bf30}(1), 13-27 (2015)
   \bibitem{MQ20}Monta¨¦s, D.C., Quiroz, A.J., Dulcerubio, M.D., Riascosvillegas, A.J.: Efficient nearest neighbors methods for support vector machines in high dimensional feature spaces. Optim. Lett. https://doi.org/10.1007/s11590-020-01616-w. (2020)

   % \bibitem{MN14} Mordukhovich, B., Nam, N.: An easy path to convex analysis and applications.  Morgan and Claypool Pubulishers. (2014)
%\bibitem{ON20} Okwuashi, O., Ndehedehe, C.E.: Deep support vector machine for hyperspectral image classification. Pattern Recognit. https://doi.org/10.1016/j.patcog.2020.107298. (2020)
%     \bibitem{PT11}Peng X. TPMSVM: A novel twin parametric-margin support vector machine for pattern recognition. \emph{Pattern Recognition}, 44(10):2678-2692. (2011)
    \bibitem{PH07}Polyak, R., Ho, S.S., Griva, I.: Support vector machine via nonlinear rescaling method. Optim. Lett. {\bf1}(4), 367-378 (2007)

    %\bibitem{RW98}Rockafellar, R.T., Wets, R.J.B.:  Variational analysis. Springer Science and Business Media. (2009)
     \bibitem{SS18}Schlkopf, B., Smola, A.J., Bach, F.: Learning with kernels: support vector machines, regularization, optimization, and beyond. MIT Press. (2018)
%\bibitem{SY19} Sharif, W., Yanto, Y., Samsudin, N.A.: An optimised support vector machine with ringed seal search algorithm for efficient text classification. Journal of Engineering Science and Technology. {\bf14}(3), 1601-1613 (2019)

         \bibitem{ST03}Shen, X.T., Tseng, G.C., Zhang, X.G., Wong, W.H.: On $\psi$-Learning. J. Am. Stat. Assoc. {\bf98}(1), 724-734 (2003)
%         \bibitem{TT99} Toh, K., Todd,  M., Tutuncu, R.:  SDPT3-a matlab software package for semidefinite programming. Optimization
%Methods and Software. {\bf11}(4), 545-581 (1999)
%          \bibitem{VP19}Venkateswarlal P, Nitta G R, Prasad A. Ensemble of texture and shape descriptors using support vector machine classification for face recognition. \emph{Journal of Ambient Intelligence and Humanized Computing}, https://doi.org/10.1007/s12652-019-01192-7. (2019)
%    \bibitem{WZ20}Wang H F, Zheng B, Yoon S W, et al. A support vector machine-based ensemble algorithm for breast cancer diagnosis. \emph{European Journal of Operational Research}, 267(2):687-699. (2018)

                 \bibitem{WL20} Wang, H.R., Lu, S.J., Zhou, Z.J.: Ramp loss for twin multi-class support vector classification. Int. J. Syst. Sci.  https://doi.org/10.1080/00207721.2020.1765047. (2020)
 \bibitem{WF17} Wang, X.M., Fan, N.,  Pardalos, P.M.: Stochastic subgradient descent method for large-scale robust chance-constrained support vector machines. Optim. Lett. {\bf11}(5), 1013-1024 (2017)
                    \bibitem{ZS09} Wang, Z., Vucetic, S.: Fast online training of ramp loss support vector machines. In: Ninth IEEE International Conference on Data Mining, pp. 569-577 (2009)
            \bibitem{YY07} Wu, Y., Liu, Y.: Robust truncated hinge loss support vector machines. J. Am. Stat. Assoc. {\bf102}, 974-983 (2007)


    \bibitem{Xu06}Xu, L., Crammer, K., Schuurmans, D.: Robust support vector machine training via convex outlier ablation. In: Proceedings of the 21st National Conference on Artificial Intelligence, pp. 536-542 (2006)
%        \bibitem{YD18}Yang L M, Dong H W. Support vector machine with truncated pinball loss and its application in pattern recognition.\emph{ Chemometrics and Intelligent Laboratory Systems}, 177:89-99. (2018)
 \bibitem{YR03}Yuille, A., Rangarajan, A.: The concave-convex procedure.  Neural Comput. {\bf15}, 915-936 (2003)

 %\bibitem{ZW20} Zhou, Z.H., Wang, W., Gang, W., Zhang, L.J.:  Introduction to the theory of machine learning.  China Machine Press. (2020)


 \bibitem{ZY08} Zhang, W., Yoshida, T., Tang, X.: Text classification based on multi-word with support vector machine. Knowledge-Based Syst. {\bf21}(8), 879-886 (2008)












%\begin{thebibliography}{99}
%\small
%   \bibitem{BM02} Bartlett P L, Mendelson S. Rademacher and Gaussian complexities: Risk bounds and structural results. \emph{Journal of Machine Learning Research}, 3:463-482. (2002)
%       \bibitem{BJ11}  Brooks J P. Support vector machines with the ramp loss and the hard margin loss. \emph{Operations Research}, 59(2):467-479. (2011)
% \bibitem{CN14} Carrizosa E, Nogales-Gmez A, Romero M  D. Heuristic approaches for support vector machines with the ramp loss. \emph{Optimization Letters}, 8(3):1125-1135. (2014)
%
%   \bibitem{CS04} Christmann A, Steinwart I. On robustness properties of convex risk minimization methods for pattern recognition. \emph{ Journal of Machine Learning Research}, 5:1007-1034. (2004)
%           \bibitem{RF06a} Collobert R, Sinz F, Weston J, et al. Trading convexity for scalability. \emph{In Proceedings of the 23rd International Conference on Machine Learning}, 201-208. (2006)
%                    \bibitem{RF06b}  Collobert R, Sinz F, Weston J, et al. Large scale transductive SVMs. \emph{Journal of Machine Learning Research}, 7:1687-1712, (2006)
% \bibitem{CV95} Cortes C, Vapnik V. Support vector networks. \emph{Machine Learning}, 20(3):273-297. (1995)
%
%       \bibitem{EB11} Ertekin S, Bottou L, Giles C L. Nonconvex online support vector machines. \emph{IEEE Transactions on Pattern Analysis and Machine Intelligence}, 33(2):368-381. (2011)
% \bibitem{HS14}Huang X L, Shi L, Suykens J A K. Ramp loss linear programming support vector machine. \emph{Journal of Machine Learning Research}, 15:2185-2211. (2014)
%%     \bibitem{HSS2014} Huang X L, Shi L, Suykens J A K. Support vector machine classifier with pinball loss. \emph{IEEE Transactions on Pattern Analysis and Machine Intelligence}, 36(5):984-997. (2014)
% \bibitem{GG19}Gaudioso M, Gorgone E, Hiriart-Urruty J B. Feature selection in SVM via polyhedral k-norm. \emph{Optimization Letters}, 14(1):19-36. (2020)
%% \bibitem{GM14} Ghosh S, Mondal S, Ghosh B. A comparative study of breast cancer detection based on SVM and MLP BPN classifier.\emph{
%%International Conference on First International Conference on Automation, Control, Energy
%%and Systems}, 87-90. (2014) %(2014)}
%% \bibitem{KK12}Khan N M, Ksantini R, Ahmad I S, et al. A novel SVM+NDA model for classification with an application to face recognition. \emph{Pattern Recognition}, 45(1):66-79. (2012)
% \bibitem{LC11}Li C H, Kuo B C, Lin C T, et al. A spatial-contextual support vector machine for remotely sensed image classification. \emph{IEEE Transactions on Geoence and Remote Sensing}, 50(3):784-799. (2012)
%      \bibitem{LS05}Liu S J, Shen X T, Wong W H. Computational developments of $\psi$-learning. \emph{In Proceedings of the 2005 SIAM International Conference on Data Mining, Society for Industrial and Applied Mathematics}. (2005)
%    \bibitem{LR20} Liu Y Y, Ren L, Cao X H, et al. Breast tumors recognition based on edge feature extraction using support vector machine. \emph{Biomedical Signal Processing and Control}, https://doi.org/10.1016/j.bspc.2019.101825. (2020)
%%\bibitem{LZ20}  Liu Z Y, Kan H P, Zhang T, et al. DUKMSVM: A framework of deep uniform kernel mapping support vector machine for short text classification. \emph{ Applied Sciences}, 10(7):1-16. (2020)
%        \bibitem{LW19} Lu S J, Wang H R, Zhou Z J. All-in-one multicategory ramp loss maximum margin of twin spheres support vector machine. \emph{Applied Intelligence}, 49(6): 2301-2314. (2019)
%    \bibitem{DE15}  Mancev D. A sequential dual method for the structured ramp loss minimization. \emph{Facta Universitatis}, 30(1):13-27. (2015)
%   \bibitem{MQ20}Monta¨¦s D C, Quiroz A J, Dulcerubio M, et al. Efficient nearest neighbors methods for support vector machines in high dimensional feature spaces. \emph{Optimization Letters}, https://doi.org/10.1007/s11590-020-01616-w. (2020)
%    \bibitem{MN14} Mordukhovich B, Nam N. An easy path to convex analysis and applications. \emph{ Morgan and Claypool Pubulishers}. (2014)
%\bibitem{ON20} Okwuashi O, Ndehedehe C E. Deep support vector machine for hyperspectral image classification. \emph{Pattern Recognition}, https://doi.org/10.1016/j.patcog.2020.107298. (2020)
%%     \bibitem{PT11}Peng X. TPMSVM: A novel twin parametric-margin support vector machine for pattern recognition. \emph{Pattern Recognition}, 44(10):2678-2692. (2011)
%    \bibitem{PH07}Polyak R, Ho S S, Griva I. Support vector machine via nonlinear rescaling method. \emph{Optimization Letters}, 1(4):367-378. (2007)
%
%    \bibitem{RW98}Rockafellar R T, Wets R J B.  Variational analysis. \emph{Springer Science and Business Media}. (2009)
%     \bibitem{SS18}Schlkopf B, Smola A J, Bach F. Learning with kernels: support vector machines, regularization, optimization, and beyond. \emph{MIT Press}, 2018.
%\bibitem{SY19} Sharif W, Yanto Y, Samsudin N A, et al. An optimised support vector machine with ringed seal search algorithm for efficient text classification. \emph{ Journal of Engineering Science and Technology}, 14(3):1601-1613. (2019)
%         \bibitem{ST03}Shen X T, Tseng G C, Zhang X G, et al. On $\psi$-Learning. \emph{Journal of the American Statal Association}, 98(1):724-734. (2003)
%         \bibitem{TT99} Toh K, Todd  M, Tutuncu R. 1999. SDPT3-a matlab software package for semidefinite programming. \emph{Optimization Methods and Software}, 11(4):545-581. (1999)
%%          \bibitem{VP19}Venkateswarlal P, Nitta G R, Prasad A. Ensemble of texture and shape descriptors using support vector machine classification for face recognition. \emph{Journal of Ambient Intelligence and Humanized Computing}, https://doi.org/10.1007/s12652-019-01192-7. (2019)
%%    \bibitem{WZ20}Wang H F, Zheng B, Yoon S W, et al. A support vector machine-based ensemble algorithm for breast cancer diagnosis. \emph{European Journal of Operational Research}, 267(2):687-699. (2018)
%
%                 \bibitem{WL20} Wang H R, Lu S J, Zhou Z J. Ramp loss for twin multi-class support vector classification. \emph{International Journal of Systems Science}, https://doi.org/10.1080/00207721.2020.1765047. (2020)
% \bibitem{WF17} Wang X M, Fan N,  Pardalos P M. Stochastic subgradient descent method for large-scale robust chance-constrained support vector machines. \emph{Optimization letters}, 11(5):1013-1024. (2017)
%                    \bibitem{ZS09} Wang Z, Vucetic S. Fast online training of ramp loss support vector machines. \emph{International Conference on Ninth IEEE International Conference on Data Mining}. (2009)
%            \bibitem{YY07} Wu Y, Liu Y. Robust truncated hinge loss support vector machines. \emph{Journal of the American Statistical Association}, 102:974-983. (2007)
%
%
%    \bibitem{Xu06}Xu L, Crammer K, Schuurmans D. Robust support vector machine training via convex outlier ablation. \emph{In Proceedings of the 21st National Conference on Artificial Intelligence}, 536-542. (2006)
%%        \bibitem{YD18}Yang L M, Dong H W. Support vector machine with truncated pinball loss and its application in pattern recognition.\emph{ Chemometrics and Intelligent Laboratory Systems}, 177:89-99. (2018)
% \bibitem{YR03}Yuille A, Rangarajan A. The concave-convex procedure.  \emph{Neural computation}, 15:915-936. (2003)
%
% \bibitem{ZW20}  Zhou Z H, Wang W, Gang W, et al.  Introduction to the Theory of Machine Learning.  \emph{China Machine Press}. (2020)
%
%
%% \bibitem{ZY08} Zhang W, Yoshida T, Tang X. Text classification based on multi-word with support vector machine. \emph{ Knowledge-Based Systems}, 21(8):879-886. (2008)






















    %\bibitem{QD19} Qiaolin Y E, Dengping X U, Dong Z. Remote sensing image classification based on deep learning features and support vector machine.\emph{ Journal of Forestry Engineering, 2019.}

%\bibitem{ZD19}Zhang F , Dong P , Wei W . Face recognition by support vector machine optimized by an improved grey wolf algorithm[J]. Computer Engineering and science, 2019.














%\bibitem{TY13} Farhan, Mohammad, Abdullah, Mujeeb, Kassem, Ghulam,. Support Vector Machine Classifier for Pattern Recognition. International Conference on Informatics & Computational Intelligence. (2012).
%    Farhan M, Kassem G, Abdullah M et al. Support vector machine classifier for pattern recognition. In 2011 First International Conference on Informatics and Computational Intelligence (pp. 272-277). IEEE.





%\bibitem{VE20}Vidi I , Egnell L , Jerome N P , et al. Support vector machine for breast cancer classification using diffusion©\weighted MRI histogram features: Preliminary study[J]. Journal of Magnetic Resonance Imaging, 2018.

%Zhu W , Bao H , Zeng Z , et al. Support Vector Machine Optimized Using the Improved Fish Swarm Optimization Algorithm and Its Application to Face Recognition[J]. International Journal of Pattern Recognition and Artificial Intelligence, 2019.










%\bibitem{GE17} Gohariyan E, Esmaeilpour M, Shirmohammadi MM. The combination of mammography and MRI for diagnosing breast cancer using fuzzy NN and SVM. \emph{International Journal of Interactive Multimedia and Artificial Intelligence, 4(5):20-24.  (2017)}

%%\bibitem{PK14}  Pang Y, Zhang K, Yuan Y et al. Distributed object detection with linear SVMs. \emph{IEEE Transactions on Cybernetics, 44(11):2122-2133.(2014)}
% \bibitem{PMY14}   Poursaeidi M H , Kundakcioglu O E . Robust support vector machines for multiple instance learning. \emph{Annals of Operations Research,     216(5):205-227. (2014)}








 %  \bibitem{LH08} Wang L, Jia H D, Li J. Training robust support vector machine with smooth Ramp loss in the primal space. \emph{Application Research of Computers,71(13):3020-3025. (2008)}
%   \bibitem{LH16}  Liu D L, Shi Y, Huang X. Ramp loss least squares support vector machine. \emph{Journal of Computational Science, 14:61-68.(2016)}
%













%\bibitem{AA02} Yullie A L, Rangarajan A, The concave-convex procedure. \emph{In Advances in Neural Information Processing Systems. (2002)}





% \bibitem{FC05} Fan, R.-E., Chen, P.-H., and Lin, C.-J., ¡°Working set selection using
%second order information for training SVM¡±, JMLR, 2005.
%\bibitem{GT08} Glasmachers, T. and Igel, C., ¡°Second order SMO improves SVM
%online and active Learning¡±, Neural Computation, 2008.
%\bibitem{Zhou2015} Zhou Z H. Maching Learning, \emph{Beijing: Tsinghua University Press}(2015)
%======================A-G==================
%\bibitem{Allen-Zhu-Hazan2016} Allen-Zhu Z, Hazan E. Variance reduction for faster non-convex optimization. {\em in Proceedings of the 33rd International Conference on Machine Learning}, 699-707.(2016)
%\bibitem{Auussel1998}Auussel D. Subdifferential properties of quasiconvex and pseudoconvex functions: Unified Approach. \emph{Journal of Optimization Theory and Applications}, 97(1): 29-45.(1998)
%\bibitem{Aussel-Corvellec1995}Aussel D, Corvellec J N , Lassonde M. Mean value property and subdifferential criteria for lower semicontinuous functions. \emph{Transactions of the American Mathematical Society}, 347(10):4147-4161.(1995)
%\bibitem{Aussel1998} Aussel D. subdifferential properties of quasivonvex and pseudoconvex functions. \emph{Journal of Optimization Theory and Applications}.(1997)
%\bibitem{Bertsekas1999} Bertsekas D P. Nonlinear programming. \emph{Belmont MA: Athena Scientific, second edition.}(1999)
%\bibitem{Beck-Eldar2013} Beck A, Eldar Y C.~Sparsity constrained nonlinear optimization: optimality conditions and algorithms. \emph{SIAM Journal on Optimization.} 23(3):1480-1509.(2013)
%\bibitem{Beck-Hallak2015} Beck A, Hallak N. On the Minimization Over Sparse Symmetric Sets: Projections, Optimality Conditions, and Algorithms. \emph{Mathematics of Operations Research}, 41(1):196-223.(2016)
%\bibitem{Bertsekas1999} Bertsekas D P. Nonlinear Programming. \emph{2nd ed,Athena Scientific}, Belmont, MA, (1999)
%%\bibitem{Beck-Hallak2018} Beck A and Hallak N. Proximal mapping for symmetric penalty and sparsity. \emph{SIAM Journal on Optimization}, 28(1): 496-527.(2018)
%%\bibitem{Bergstra-Desjardins2009}Bergstra J, Desjardins G, Lamblin P, Bengio Y. Quadratic polynomials learn better image features. \emph{Technical report}, 1337. (2009)
%\bibitem{Bian2019}Bian W. A smoothing proximal gradient algorithm for nonsmooth convex regression with cardinality penalty. Submitted: 1-25.(2019)
%%\bibitem{Boris-Nguyen2014} Boris S M,~Nguyen M N.~An easy path to convex analysis and application. \emph{Synthesis Lectures on Mathematics and Statistics}, (2014)
%%\bibitem{Borwein-Lewis2000} Borwein J, Lewis A S. Convex analysis and nonlinear optimization: theory and examples. \emph{Cms Books in Mathematics}. (2000)
%\bibitem{Blumensath2013}Blumensath T.~Compressed sensing with nonlinear observations and related nonlinear optimization problems. \emph{IEEE Transactions on Information Theory}, 59(6):3466-3474.(2013)
%%\bibitem{Belloni-Chernozhukov2011} Belloni A, Chernozhukov V. L1-penalized quantile regression in high-dimensional sparse models,\emph{The Annals of Statistics}, 39: 82-130 (2011)
%%\bibitem{Bradic-Fan2011}Bradic J, Fan J, Wang W. Penalized composite quasi-likelihood for ultrahigh-dimensional variable selection. \emph{Journal of the Royal Statistical Society}, 73(3):325-349.(2009)
%\bibitem{Blumensath-Davies2008}Blumensath T, Davies M E. Iterative thresholding for sparse approximations. \emph{Journal of Fourier Analysis and Applications}, 14(5-6):629-654.(2008)
%\bibitem{Blumensath-Davies2009} Blumensath T, Davies M E. Iterative hard thresholding for compressed sensing. \emph{Applied and Computational Harmonic Analysis}, 27(3):265-274.(2009)
%%\bibitem{Bucher-Schwartz2018} Max B, Alexandra S. Second-Order Optimality Conditions and Improved Convergence Results for Regularization Methods for Cardinality-Constrained Optimization Problems. \emph{Journal of Optimization Theory and Applications}, 178(2):383-410.(2018)
%\bibitem{Chen-Xiu2013} Chen Y Q, Xiu N H, Peng D T. Global solutions of non-Lipschitz $S_2¨CS_p$ minimization over the positive semidefinite cone. Optimization Letters, 8(7):2053-2064.(2013)
%%\bibitem{Chen-Pan2017}Chen X J, Pan L L, Xiu N H.~Relationship between three sparse optimization problems for multivariate regression.(2017)
%{\color{blue}\bibitem{Chen-Pan2019} Chen X J, Pan L L, Xiu N H. Relationship between three sparse optimization problems for multivariate regression. Submitted: 1-32.(2019)}
%\bibitem{Chib1992} Chib S. Bayes inference in the Tobit censored regression model. \emph{Journal of Econometrics}, 51(1-2):79-99.(1992)
%\bibitem{Clarke1983} Clarke F H. Optimization and nonsmooth analysis,\emph{ Wiley}.(1983)
%{\color{blue}\bibitem{Clarke1989} Clarke F H. Methods of dynamic and nonsmooth optimization, \emph{CBMS-NSF Regional conference series in Applied mathmatics, SIAM Publications, Philadelphia}, 57. (1989)}
%%\bibitem{Clevert-Unterthiner2016} Clevert D A, Unterthiner T. Sepp hochreiter fast and accurate deep network learning by exponential linear units(ELUs). \emph{International Conference on Learning Representations},(2016)
%%\bibitem{Cartis-Gould2011} Cartis C, Gould N I M, Toint P L. Adaptive cubic regularisation methods for unconstrained optimization. Part I: motivation, convergence and numerical results.\emph{ Mathematical Programming}, 127(2):245-295.(2011)
%
%%\bibitem{Candes2006} Cand$\grave{e}$s E J. Compressive sampling. Proceedings of the International congress of mathematicians, Switzerland: European Mathematical Society Publishing House, 1433-1452. (2006)
%
%\bibitem{Candes-Tao2005} Cand$\grave{e}$s E J, Tao T. Decoding by linear programming. \emph{IEEE Transactions on Information Theory}, %51(42):4203-4215.(2005)
%\bibitem{Chen-Ge2014} Chen X J, Ge D D, Wang Z Z, et al. Complexity of unconstrained $L_2-L_p$ minimization. \emph{Mathematical Programming}, 143(1-2):371-383.(2014)
%\bibitem{Cui-Pang2018} Cui Y , Pang J S , Sen B. Composite Difference-Max Programs for Modern Statistical Estimation Problems. \emph{SIAM Journal on Optimization}, 28(4):3344-3374. (2018)
%%\bibitem{Chen-Guo2017} Chen X J, Guo L, Lu Z S, et al. An Augmented Lagrangian Method for Non-Lipschitz Nonconvex Programming. \emph{SIAM Journal on Numerical Analysis}, 55(1):168-193.(2017)
%%Chouzenoux,E. Jezierska A., Pesquet J.-C. , and H. Talbot, A majorize-minimize subspace approach for $l_2 -l_0$ image regularization, SIAM J. Imaging Sci., 6, 563-591(2013).
%\bibitem{Donoho2006} Donoho D L. Compressed sensing. \emph {IEEE Transactions on Information Theory}, 52(4):1289-1306.(2006)
%%\bibitem{Fan-Fan2014}Fan, Y., Fan, J., Barut, E.: Adaptive robust variable selection. Ann. Stat. 42(1), 324-351 (2014)
%%\bibitem{Fan-Li2001}Fan J Q, Li R Z. Variable selection via nonconvave penalized likelihood and its oracle properties,\emph{ Journal of the American Statistical Association}, 1348-1360 (2001)
%%\bibitem{He-Zhang2015} He K M, Zhang X Y, Ren S Q, Sun J. Delving Deep into Rectifiers: Surpassing Human-Level Performance on ImageNet Classification. \emph{IEEE International Conference on Computer Vision} (2015)
%%\bibitem{Ge-Jiang2011} Ge D D, Jiang X Y, Ye Y Y. A note on the complexity of $L_p$ minimization. \emph{Technical report
%\bibitem{Guo-Ye2017} Guo L, Ye J J. Necessary optimality conditions and exact penalization for non-Lipschitz nonlinear programs. \emph{Mathematical Programming}, 168(1-2):571¨C598.(2018)
%\bibitem{Glorot-Bordes2011} Glorot X, Bordes A, Bengio Y. Deep sparse rectifier neural networks. \emph{Artificial Intelligence and Statistics}. 15:315-323.(2011)
%
%%======================H-N==================
%%\bibitem{Han-Moraga2005} Han J., Moraga C. The influence of the sigmoid function parameters on the speed of backpropagation learning. Springer, Berlin, Heidelberg(2005)
%\bibitem{Hinton2010} Hinton G E. Rectified Linear Units Improve Restricted Boltzmann Machines Vinod Nair. \emph{International Conference on International Conference on Machine Learning. Omnipress.} (2010)
%\bibitem{Hossein-Ajmal2016} Hossein R, Ajmal M, Mubarak S. Learning a deep model for human action recognition from novel viewpoints. \emph{IEEE Transactions on Pattern Analysis and Machine Intelligence}, 40(3):667-681.(2017)
%
%%\bibitem{Hiriart-Urruty-Le2014} Hiriart-Urruty J B,  Le H. The Viscosity Subdifferential of the Rank Function via the Corresponding Subdifferential of its Moreau Envelopes, \emph{Acta Mathematica Vietnamica}, 40(4): 735-746(2015)% ´Î΢·ÖÖ®¼äµÄ¹Øϵ
%\bibitem{Kyunghyun-Bart2014} Cho K, Van Merrienboer B, Gulcehre C, et al. Learning Phrase Representations using RNN Encoder-Decoder for Statistical Machine Translation. \emph{Computer Science}.(2014)
%\bibitem{Le2013}Le H Y, Generalized subdifferentials of the rank function.~\emph{Optimization Letters},~7(4):731-743.(2013)
%%\bibitem{Li-Song2017}Li X R, Song W, Xiu N H.Optimality conditions for Rank-constrained Matrix Optimization,2017.
%%\bibitem{Li-Song2018}Li X R, Xiu N H, Zhou S L. Optimality conditions for Rank-constrained Matrix Optimization,2018.
%%\bibitem{Le-Thi2018} Le Thi H A, Pham Dinh T. DC programming and DCA: thirty years of developments, Mathematical Programming, Special Issue: DC Programming-Theory,~\emph{Algorithms and Applications}, 169:5-64.(2018)
%%\bibitem{Lai-Xu-Yin2013} Lai M, Xu Y, and Yin W. Improved iteratively reweighted least squares for unconstrained smoothed $l_q$ minimization. \emph{SIAM Journal on Numerical Analysis}, 51(2):927-957.(2013)
%\bibitem{Liu-Cosman2017} Liu J, Cosman P C, Rao B D. Robust linear regression via $l_0$ regularization. \emph{IEEE Transactions on Signal Processing}, 66(3):698-713.(2017)
%\bibitem{Lu2013}Lu Z S, Zhang Y. Sparse Approximation via Penalty Decomposition Methods. \emph{SIAM Journal on Optimization}, 23(4):2448-2478.(2013)
%\bibitem{Lu2014} Lu Z S. Iterative reweighted minimization methods for $l_p$-regularized unconstrained nonlinear programming. \emph{Mathematical Programming}, 147(1-2):277-307.(2014)
%\bibitem{LeCun-Bengio2015} LeCun Y, Bengio Y, Hinton G. Deep learning. \emph{Nature}, 521:436-444.(2015)
%%\bibitem{Mukkamala-Ochs2019} Mukkamala M C, Ochs P, Pock T, et al. Convex-Concave backtracking for inertial bregman proximal gradient algorithms in non-Convex optimization. \emph{arXiv}:1904.03537.(2019)
%%\bibitem{Mordukhovich-Shao1995} Mordukhovich B S , Shao Y. On nonconvex subdifferential calculus in Banach spaces. Journal of Convex Analysis, (1-2):~211-227.(1995)
%\bibitem{Mordukhovich2018} Mordukhovich B S. Variational analysis and application. \emph{Springer}.(2018)
%%\bibitem{Movahedian-Nobakhtian2017} Movahedian N, Nobakhtian S. Sarabadan M. Nonsmooth sparsity constrained optimization problems:optimality conditions. \emph{Optimization Letters}.(2017)
%%\bibitem{Marko-Pekka1992} Marko M M, Pekka N. Nonsmooth Optimization: Analysis and Algorithms with Applications to Optimal Control, \emph{World Scientific}, 1992.
%\bibitem{Natarajan1995} Natarajan B K . Sparse Approximate Solutions to Linear Systems. \emph{SIAM Journal on Computing}, 24(2):227-234.(1995)
%{\color{blue}\bibitem{Nikolova2016} Nikolova M. Relationship between the optimal solutions of least squares regularized with $l_0$-norm and constrained by k-sparsity. \emph{Applied and Computational Harmonic Analysis}, 41(1):237-265.(2016)}
%%======================O-T==================
%
%%\bibitem{Olshausen-Field1996} Olshausen B A, Field D J. Emergence of simple-cell receptive field properties by learning a sparse code for natural images, \emph{ Nature}, 381 607-609.(1996)
%%\bibitem{Ochs-Fadili2018} Ochs P, Fadili M J, Brox T. Non-smooth Non-convex bregman minimization: unification and new algorithms, Journal of Optimization Theory and Applications.(2018)
%%\bibitem{Pan-Xiu2015} Pan L L, Xiu N H, Zhou S L. On Solutions of Sparsity Constrained Optimization. \emph{Journal of the Operations Research Society of China}, 3(4):421-439.(2015)
%% \bibitem{Pan-Xiu2017} Pan L L, Xiu N H, Fan J.~Optimality conditions for sparse nonlinear programming. \emph{Science China(Mathematics)}, (05):5-22.(2017)
%\bibitem{Powell1984}Powell J L. Least absolute deviations estimation for the censored regression model. \emph{Journal of Econometrics}, 25(3):303-325.(1984)
%%
%%\bibitem{Rockafellar-Wets1998} Rockafellar R T. Wets R J.~Variational analysis. \emph{Springer}.(1998)
%%\bibitem{Rockafellar1970} Rockafellar R T.~Convex analysis. \emph{Princeton University Press}.(1970)
%%\bibitem{Thorarinsdottir2010} Thorarinsdottir T L, Gneiting T. Probabilistic forecasts of wind speed: ensemble model output statistics by using heteroscedastic censored regression. \emph{Journal of the Royal Statistical Society Series A (Statistics in Society)}, 173(2):371-388.(2010)
%%\bibitem{Witten-Frank2000} Witten I H, Frank E. Data mining: practical machine learning tools and techniques with java implementations. \emph{Morgan Kaufmann Publishers}.(2000)
%%\bibitem{Yu-Deng2015} Yu D, Deng L. Automatic Speech Recognition: A Deep Learning Approach, Signals and Communications Technology, \emph{Springer}. (2015)
%{\color{blue}
%\bibitem{Yuan-Liu2014} Yuan X T, Liu Q S. Newton greedy pursuit: A quadratic approximation method for sparsity-constrained optimization. \emph{In Proceedings of the IEEE Conference on Computer Vision and Pattern Recognition}, 4122-4129. (2014)
%\bibitem{Yuan-Liu2017} Yuan X T, Liu Q S. Newton-type greedy selection methods for $l_0$-constrained minimization. \emph{IEEE transactions on pattern analysis and machine intelligence}, 39(12): 2437-2450.(2017)
%\bibitem{Wang2019} Wang R, Xiu N H, Zhou S L. Fast newton method for sparse logistic regression. \emph{arXiv preprint arXiv:1901.02768}.(2019)
%\bibitem{Zhou2019} Zhou S L, Xiu N H, Qi H D. Global and quadratic convergence of Newton hard-thresholding pursuit. \emph{arXiv preprint arXiv:1901.02763}.(2019)
%\bibitem{Zhang-Li2017} Zhang N, Li Q. On optimal solutions of the constrained $l_0$ regularization and its penalty problem. \emph{Inverse Problems}, 33(2).(2017)}
%%\bibitem{Xiu-Kong2018}Xiu X C, Kong L C, Li Y, Qi HD.,:~Interative reweighted methods for $l_1-l_p$ minimization.2018
%%\bibitem{Zeng-Lin-Wang2014} Zeng J, Lin S, Wang Y, et al. $L_{1/2}$ Regularization: convergence of iterative half thresholding algorithm. \emph{IEEE Transactions on Signal Processing}, 62(9):2317-2329.(2014)




\end{thebibliography}

 \UseRawInputEncoding 
\RequirePackage{fix-cm}

\documentclass[smallextended]{svjour3}       
\smartqed  
\usepackage[square, comma, sort&compress, numbers]{natbib}
\usepackage{graphicx}
\usepackage[misc]{ifsym}
\usepackage{amssymb}
\usepackage{mathrsfs}
\usepackage{bm}
\usepackage{float}  
\usepackage{subfigure}  

\usepackage{amsmath}
\allowdisplaybreaks[4]
\usepackage{array}
\usepackage{graphicx}
\usepackage{cases}
\usepackage{color}
\usepackage{multirow}
\usepackage{picinpar}
\graphicspath{ {./images/} }

\usepackage[all]{xy}
\usepackage[colorlinks,
            linkcolor=blue,
            anchorcolor=green,
            citecolor=blue
            ]{hyperref}

\def\N{\mathcal{N}}

\def\N{{\mathbb N}}

\def\R{{\mathbb R}}

\def\bfe{{\bf 1}}
\def\bfe{{\bf 1}}

\def\bfe{{\bf 1}}

\def\bfv{{\bf v}}
\def\bfx{{\bf x}}
\def\bfw{{\bf w}}
\def\bfy{{\bf y}}

\def\bfs{{\bf s}}
\def\bfu{{\bf u}}

\def\R{\mathbb{R}}

\newcommand{\ba}{\begin{array}}
\newcommand{\ea}{\end{array}}
\newcommand{\be}{\begin{equation}}
\newcommand{\ee}{\end{equation}}

\newcommand{\RNum}[1]{\lowercase\expandafter{\romannumeral #1\relax}}

\begin{document}
\title{Proximal Operator and Optimality Conditions for Ramp Loss SVM
}

\titlerunning{POOC for $L_r$-SVM }        
\author{Huajun Wang$^1$ \and
        Yuanhai Shao$^2$  \and
        Naihua Xiu$^1$
}

\institute{Huajun Wang \at huajunwang@bjtu.edu.cn\and
           Yuanhai Shao \at shaoyuanhai@hainanu.edu.cn\and
           Naihua Xiu \at nhxiu@bjtu.edu.cn\and
           1. Department of Mathematics, School of Science,
Beijing Jiaotong University, Beijing \at 100044, People's Republic of China\and
 2. School of Management, Hainan University, Haikou, \at 570000, People's Republic of China
           }

\date{Received: date / Accepted: date}

\maketitle

\begin{abstract}

Support vector machines with ramp loss (dubbed as $L_r$-SVM) have attracted wide attention due to the boundedness of ramp loss. However, the corresponding optimization problem is non-convex and the given Karush-Kuhn-Tucker (KKT) conditions are only the necessary  conditions. To enrich the optimality theory of $L_r$-SVM and go deep into its statistical nature, we first introduce and analyze the proximal operator for ramp loss, and then establish a stronger optimality conditions: P-stationarity, which is proved to be the first-order necessary and sufficient conditions for local minimizer of $L_r$-SVM. Finally, we define the $L_r$ support vectors based on the concept of P-stationary point, and show that all $L_r$ support vectors fall into the support hyperplanes, which possesses the same feature as the one of hard margin SVM.

\keywords{ Ramp loss SVM \and  proximal operator \and minimizer \and P-stationary point \and  support vectors}
\end{abstract}
\vskip 2mm
\section{Introduction}
 \vskip 3mm

Support vector machines (SVM) were first introduced by Vapnik and Cortes \cite{CV95} and have been  widely
 applied into many fields, including text and image classification \cite{ZY08,LZ20}, disease detection \cite{GG19,PH07}, etc.
The decision hyperplane of SVM classifier, $\langle  \bfw , \bfx\rangle+b=0$ with $ \bfw \in \R^n$ and $b\in \R$, is trained from data set $\{(\bfx_i,y_i), i\in\N_m\}$ where $\bfx_i\in \R^n$, $y_i\in \{-1,1\}$ and $\N_m:=\{1,2,\cdots,m\}$ by optimizing
the following problem
\begin{eqnarray}\label{SM-SVM}
\min_{ \bfw \in \R^n,b\in \R}~~\frac{1}{2}\Vert \bfw \Vert^{2}+C\sum_{i=1}^{m}\ell_h(1-y_{i}( \langle \bfw, \bfx_i\rangle +b)),
\end{eqnarray}
where $C>0$ is a penalty parameter and $\ell_h(t)=\max\{t,0\}$ is the hinge  loss, which has no cost for $t<0$, but pays linear cost for $t\geq0$. The cost given to the outliers by the hinge loss is quite huge since it is unbounded function \cite{HS14,DE15,MQ20,WF17,KS09,LS05}.  To remove the impact of outliers, one method for increasing the robustness of SVM is to use the ramp loss \cite{ST03} (see Fig. 1 (a)).  The ramp loss is defined as follows,
\begin{eqnarray}\label{Rloss}
\ell_{r}(t)=
\begin{cases}
1,& t\geq1,\\
t,& 0\leq t<1,\\
0,& t<0,
\end{cases}
\end{eqnarray}
which is also called truncated hinge loss in \cite{YY07}. It has no cost for $t<0$, but it pays linear cost for $0\leq t< 1$ and a fixed cost at 1 for $t\geq1$. Authors in \cite{RF06a,RF06b} also extended and studied ramp loss by replacing all constant 1 in (\ref{Rloss}) by adjustable parameter $\mu$ $(\mu>0)$.  Since the theoretical results have no essential difference between the fixed 1 and parameter $\mu$ for ramp loss, for simplicity of the proof, we use formula \eqref{Rloss} directly. Bartlett et al. \cite{BM02} investigated some of the theoretical
properties of the ramp loss, while SVM with ramp loss  (dubbed as $L_r$-SVM) was first proposed by Shen et al. \cite{ST03}. Replacing $\ell_{h}(\cdot)$ by $\ell_{r}(\cdot)$ in (\ref{SM-SVM}), the optimization problem of $L_r$-SVM is 
\begin{eqnarray} \label{R-SVM}
\min_{ \bfw \in {\R}^n, b\in {\R}} f_r( \bfw ;b):=\frac{1}{2}\Vert \bfw  \Vert^2+CL_r({\bfe}-A \bfw -b{\bf{y}}),
\end{eqnarray}
where $A:=[y_1\bfx_{1}~y_2\bfx_{2}~\cdots~y_m\bfx_{m}]^\top\in {\R}^{m\times n}, \bfy:=(y_{1},y_2,\cdots,y_{m})^{\top}\in {\R}^{m}, {\bfe} :=(1,1,\cdots,1)^{\top}\in {\R}^{m}$, $L_r(\bfu):=\sum_{i=1}^m \ell_{r}(u_i)$ with $\bfu:=(u_1,u_2,\cdots,u_m)^\top={\bfe}-A \bfw -b{\bf{y}}\in\R^m$, which computes the sum of positive elements for all $u_i<1$ and the number of elements for all $u_i\geq 1$ in $\bfu$. Due to the non-convexity of ramp loss, the $L_r$-SVM (\ref{R-SVM}) is a non-differentiable non-convex optimization problem. Thus, to find a satisfying solution, scholars have paid many efforts in optimality conditions and algorithms for $L_r$-SVM.

 In recent decades, on the one hand, some approaches such as convex relaxation and equivalent reformulation as mixed integer nonlinear programming (MINP) problem were proposed for dealing with $L_r$-SVM. Xu et al. \cite{Xu06} reformulated $L_r$-SVM as a semidefinite programming problem by convex relaxation which was solved by the MATLAB software package SDPT3.  Brooks \cite{BJ11}  transformed  $L_r$-SVM into a MINP problem, and then proposed a branch and bound algorithm to solve it. Carrizos et al. \cite{CN14} developed  heuristic method to handle the  MINP problem of $L_r$-SVM on large datasets.

On the other hand, Collobert et al. \cite{RF06a} translated $L_r$-SVM (\ref{R-SVM}) into an equivalent  difference of convex functions (DC) programming and took the subdifferential of $f_r(\bfw;b)$ to establish the first-order necessary conditions: Karush-Kuhn-Tucker (KKT) conditions, and used the concave-convex
procedure (CCCP) \cite{YR03} for solving the DC programming.  To improve the computational speed, Wang et al. \cite{ZS09} designed an efficient working set selection strategy based on KKT conditions and then used the CCCP with working set for solving the DC programming.
 Since the KKT conditions provided effective characterization of optimal solution to $L_r$-SVM (\ref{R-SVM}), this class of approaches continued to be widely studied in theory as well as algorithms. For more details, see, e.g., \cite{EB11,WL20,LW19,SS18} and references therein.

 A natural question arises. Is there stronger first-order optimality conditions of $L_r$-SVM (\ref{R-SVM}) to provide more effective characterization of optimal solution?
It is this question that motivates the work in our paper.
 The main results of this paper are summarized as follow: (i) We derive the explicit expression of proximal operator of ramp  loss.  (ii) We introduce a novel optimality conditions: P-stationarity, which is proved to be the necessary and sufficient conditions for local minimizer of $L_r$-SVM. (iii) We prove that the set of the support vectors defined by P-stationary point fall into the support hyperplanes.


This paper is organized as follows. In Section 2, we derive the explicit expression of proximal operator of ramp  loss. In Section 3, we introduce a concept of P-stationary point, and reveal the relationship between P-stationary point and local/global minimizer of $L_r$-SVM. In Section 4, we introduce $L_r$ support vectors based on the P-stationary point and discuss its properties. Conclusions are made in Section 5.
\vskip 1mm
\section{Proximal operator for ramp loss}
In this section,  we derive the explicit expression of proximal operator of ramp loss, which will be used to study the new first-order optimality conditions of $L_r$-SVM in the next section.
{\bf \subsection{ $\ell_{r}$  proximal operator }}
We first give the definition of $\ell_{r}$ proximal operator in one-dimensional case.
\begin{definition}
($\ell_{r}$ proximal operator) For any given $\gamma, C>0$ and ${s}\in \R$, the proximal operator of $\ell_{r}(v)$ (dubbed as $\ell_{r}$ proximal operator) is defined as
\begin{eqnarray} \label{proxmalL}
\text{prox}_{\gamma C \ell_{r}}({ s})=\arg \min_{v\in {\R}}~ C
\ell_{r}(v)+\frac{1}{2\gamma}(v-s)^2.
\end{eqnarray}\end{definition}

The following two propositions state that the $\ell_{r}$ proximal operator admits a closed form solution for  $0<\gamma C<2$ or $\gamma C\geq2$.
\begin{proposition}\label{pro1} {\rm(Solution to $\ell_{r}$ proximal operator for $0<\gamma C<2$) \label{proximalLma1} For any given $\gamma, C>0$ and $0<\gamma C<2$, the solution to $\ell_{r}$ proximal operator  at $s \in \R$ is given as
}\end{proposition}
\be \label{prox1}
\text{prox}_{\gamma C\ell_{r}}(s):=
\begin{cases}
s,& s>1+\frac{\gamma C}{2},\\
s ~\text{or}~ s-\gamma C,& s=1+\frac{\gamma C}{2},\\
{s-\gamma C} ,&\gamma C\leq s<1+\frac{\gamma C}{2},\\
{0} ,&0< s< \gamma C,\\
s,&  \ s\leq0.
\end{cases}
\ee

\begin{proposition}{ \label{prox2}\rm(Solution to $\ell_{r}$ proximal operator for $\gamma C\geq2$) \label{proximalLma1} For any given $\gamma, C>0$ and $\gamma C\geq2$, the solution to $\ell_{r}$ proximal operator  at $s \in \R$ is given as
}\end{proposition}
\be \label{prox12}
\text{prox}_{\gamma C\ell_{r}}(s):=
\begin{cases}
s,& s>\sqrt{2\gamma C},\\
 s~\text{or}~0,& s=\sqrt{2\gamma C},\\
0 ,&0< s<\sqrt{2\gamma C},\\
s,&  \ s\leq0.
\end{cases}
\ee

{\bf\subsection{ $L_{r}$ proximal operator }}
Based on separate property of $L_{r}(\cdot)$, we extend \eqref{prox1} and \eqref{prox12} to  multi-dimensional case.
\vskip 2mm
\begin{definition}($L_{r}$ proximal operator) \label{peox_define}For any given $\gamma, C>0$, the proximal operator of $L_{r}(\bfv)$ (dubbed as $L_{r}$ proximal operator) at ${\bf s}=(s_1,s_2,\cdots,s_m)^{\top}\in\R^m$ is defined as
\begin{eqnarray} \label{proxmalLL}
\text{prox}_{\gamma CL_{r}}({\bf s})=\arg\min_{\bfv\in {\R}^m}~ C
L_{r}(\bfv)+\frac{1}{2 \gamma}\|\bfv-{\bf s}\|^2.
\end{eqnarray}
\end{definition}

The following proposition states that the $L_{r}$ proximal operator admits a closed form solution for two cases: $0<\gamma C<2$ or $\gamma C\geq2$.
\begin{proposition}\label{pro3}{\rm(Solution to $L_{r}$ proximal operator) \label{multi}For a given $\gamma, C>0$, the solution to $L_{r}$ proximal operator at ${\bf s}=(s_1,s_2,\cdots,s_m)^{\top}\in\R^m$ is given as
\begin{eqnarray}\label{Prox_L}
\text{prox}_{\gamma CL_{r}}({ \bfs}):=
\left[\begin{array}{c}
\text{prox}_{\gamma C\ell_{r}}(s_1)\\
\vdots\\
{ \text{prox}_{\gamma C\ell_{r}}(s_m)}
\end{array}\right],\end{eqnarray}
where $\text{prox}_{\gamma C\ell_{r}}(s_i)$ takes formula (\ref{prox1}) as $0<\gamma C<2$ or (\ref{prox12}) as $\gamma C\geq2$.
}\end{proposition}

\noindent {\bf \emph{Proof}} It follows from (\ref{proxmalLL}) that $[\text{prox}_{\gamma CL_{r}}({ \bfs})]_i=\text{prox}_{\gamma C\ell_{r}}({ s_i})$, where
\begin{align*}
\text{prox}_{\gamma C\ell_{r}}(s_i)=\text{arg}\min_{v\in {\R}} C\ell_{r}(v)+ \frac{1}{2\gamma}(v-s_i)^{2}, i\in \N_m.
\end{align*}
 Using (\ref{prox1}) or (\ref{prox12}) completes the proof.  \hfill  $ \Box $\vskip 2mm

\section{ First-order optimality conditions}

In this section, we develop a first-order necessary and sufficient optimality conditions for (\ref{R-SVM}). To proceed this, we introduce a variable ${\bfu}\in {\R}^{m}$ to equivalently  reformulate (\ref{R-SVM}) as
\begin{eqnarray} \label{RR-SVM}
&\underset{ \bfw \in {\R}^n,b \in \R,{\bfu} \in {\R}^m}{\min} & \frac{1}{2} \Vert \bfw  \Vert^2+CL_r(\bfu)\\ \nonumber
&\mbox{s.t.} &{\bfu}+A  \bfw +b\bfy={\bfe}.
\end{eqnarray}
Now let us define some notation
\begin{eqnarray}\label{BH}
B:=[A~\bfy]\in\R^{m\times(n+1)},~~~~
H:=\begin{bmatrix}
I_{n\times n} & {\bf0}  \\
{\bf0} & 0
\end{bmatrix}B^\dag,
\end{eqnarray}
 where  $B^\dag\in\R^{(n+1)\times m}$ is the generalized inverse of $B$ and $I_{n\times n}$ is the identity matrix with order $n$. Denote
$\lambda_H:= \lambda_{\max}(H^\top H)$ where $\lambda_{\max}(H^\top H)$ is the maximum eigenvalue of $H^\top H.$
\begin{definition}
(P-stationary point of (\ref{RR-SVM})) For a  given $C>0$, we call $({ \bfw^* }; b^*;  \bfu^* )$ is a proximal stationary (P-stationary) point of (\ref{RR-SVM}) if  there exists a Lagrangian multiplier vector $ {\bm{\lambda}^*}\in \R^m$ and a constant $\gamma>0$ such that
\be\label{PH-sta}
\left\{
\begin{array}{rll}   \bfw^* + A^{\top}{\bm{\lambda}^*} &=& {\bf 0},
\\ \langle \bfy, {\bm{\lambda}^*} \rangle&=&  0,
\\  \bfu^*+A   \bfw^*+b^*\bfy&=&
{\bfe},
\\ \text{prox}_{\gamma C L_{r}}(\bfu^*-\gamma{\bm{\lambda}^*})&\ni&\bfu^*.
\end{array}\right.
\ee
\end{definition}

Based on the above definition, we obtain the desired result in this section.

\begin{theorem}{\rm\label{theom1} (First-order necessary optimality conditions). Let $B$ be full  column rank. For a given $C>0$, if $( \bfw ^*; b^*; {\bfu}^*)$ is a global minimizer of (\ref{RR-SVM}), then it is a P-stationary point with $0<\gamma< 1/\lambda_H$.

 }\end{theorem}

 \begin{theorem}{\rm\label{theom2} (First-order sufficient optimality conditions). For a given $C>0$, if $( \bfw ^*; b^*; {\bfu}^*)$ with ${\bm{\lambda}}^*\in {\R}^m$ and $\gamma>0$ is a P-stationary point, then it is a local minimizer of (\ref{RR-SVM}).
 }\end{theorem}

\vskip 2mm
At the end of this section, we analyze the  following relationship between the P-stationary point  and the KKT point given by \cite{RF06a}. Based on the problem of \eqref{RR-SVM} and the necessary conditions results in \cite{RF06a}, if $( {\bf\overline{w}} ; \overline{b}; {\bf\overline{u}} )$ is a local minimizer of (\ref{RR-SVM}), then we have the following KKT conditions,

\be\label{PH-stakkt}
\left\{
\begin{array}{rll}   \overline{\bfw} + A^{\top}{\bm{\overline{\lambda}}} &=& {\bf 0},
\\ \langle \bfy, {\bm{\overline{\lambda}}} \rangle&=&  0,
\\  \overline{\bfu}+A   \overline{\bfw}+\overline{b}\bfy&=&
{\bfe},
\\ C\partial L_{r}(\overline{\bfu})+\bm{\overline{\lambda}}&\ni&{\bf 0},
\end{array}\right.
\ee
where ${\bm{\overline{\lambda}}} \in {\R}^m$ is a multiplier vector, $( {\bf\overline{w}} ; \overline{b}; {\bf\overline{u}} )$ is called the KKT point of \eqref{RR-SVM}, and $\partial L_r({\bf\overline{u}})=(\partial \ell_r(\overline{u}_1),\partial \ell_r(\overline{u}_2),\cdots,\partial \ell_r(\overline{u}_m))^\top\in \R^m$ is the  subdifferential of $L_r({\bf\overline{u}})$ with 
\be\nonumber
\partial\ell_r(\overline{u}_i)
\begin{cases}
\in[0,1],& \overline{u}_i\in\{0,1\},\\
=1,& \overline{u}_i\in(0,1),\\
=0,&\overline{u}_i<0~\text{or}~\overline{u}_i>1 ,
\end{cases}~~i\in \N_m.
\ee
Furthermore, from the above formula and last formula of \eqref{PH-stakkt}, we have
\begin{equation}\label{KKT-lam}
\textbf{0}\in C\partial
L_r({\bf\overline{u}})+{\bm{\overline{\lambda}}} \Leftrightarrow \left\{\bm{\overline{\lambda}}\in {\R}^m:
\overline{\lambda}_i
\begin{cases}
 \in[-C,0], &\overline{u}_i\in\{0,1\},\\
=-C,&\overline{u}_i\in(0,1),
\\ =0,&\overline{u}_i<0~\text{or}~\overline{u}_i>1.
 \end{cases}
 \right.
\end{equation}


\begin{theorem}{\rm\label{P-KKT}  For a given $C>0$, if $( \bfw ^*; b^*; {\bfu}^* )$ with $\bm{{\lambda}}^*\in\R^m$ and $\gamma>0$ is a P-stationary point of (\ref{RR-SVM}), then it is also a KKT point of (\ref{RR-SVM}),  but the converse does not hold.
}\end{theorem}
\noindent {\bf \emph{Proof}} The former conclusion  is obvious since the P-stationary point is  the local minimizer of (\ref{RR-SVM}). The latter conclusion is from the following counterexample. Consider the training set with the positive vectors $\bfx_1=(3,3)^\top$, $\bfx_2=(6,-2)^\top$ and negative
vector $\bfx_3=(1,1)^\top$. Namely,
\begin{eqnarray*}
A:=[y_1\bfx_{1}~y_2\bfx_{2}~y_3\bfx_{3}]^\top=
\begin{bmatrix}
3&6&-1\\
3&-2&-1
\end{bmatrix}^\top,~~\bfy=\begin{bmatrix}
 1,&1,&-1
\end{bmatrix}^\top.
\end{eqnarray*}
For a given $C=0.25$, we can verify that
\begin{eqnarray*}
{\bf\overline{w}}&=&(0.5,0.5)^\top,~~~~~~~~~~~~~~~ \overline{b}=-2,~~~~~~~~~ {\bf\overline{u}}=(0,1,0)^\top,\\
\bm{\overline{\lambda}}&=&(-0.25,0,-0.25)^\top, ~~~~\partial
L_r({\bf\overline{u}})=(1,0,1)^\top
\end{eqnarray*}
 satisfy \eqref{PH-stakkt}. This means  $( {\bf\overline{w}} ; \overline{b}; {\bf\overline{u}} )$ with $\bm{\overline{\lambda}}$ is a KKT point of \eqref{RR-SVM}.  Particularly, for $i=2$, we have $\overline{u}_2=1$ and $\overline{\lambda}_2=0$. However, for $0<\gamma<8$ and $C=0.25$, i.e., $0<\gamma C<2$, from \eqref{prox1} and $\overline{u}_2-\gamma\overline{\lambda}_2\in (0, 1+\frac{\gamma C}{2})$, we obtain $\text{prox}_{\gamma C\ell_{r}}(\overline{u}_2-\gamma\overline{\lambda}_2)=0~ \text{or}~ \overline{u}_2-\gamma C$ but the both does not equal $\overline{u}_2$. For $\gamma\geq8$ and $C=0.25$, i.e., $\gamma C\geq2$, from \eqref{prox12} and $\overline{u}_2-\gamma\overline{\lambda}_2\in (0, \sqrt{2\gamma C})$, we get $\text{prox}_{\gamma C\ell_{r}}(\overline{u}_2-\gamma\overline{\lambda}_2)= 0\neq \overline{u}_2$. To sum up, we have $\overline{u}_2\notin\text{prox}_{\gamma C\ell_{r}}(\overline{u}_2-\gamma\overline{\lambda}_2)$, which means $( {\bf\overline{w}} ; \overline{b}; {\bf\overline{u}} )$  is not a P-stationary point of (\ref{RR-SVM}). \hfill $\Box $

\section{$L_r$ support vectors}\label{sec:l01-sv}

In this section, we define the support vectors  for $L_r$-SVM by P-stationary point of (\ref{RR-SVM}), which are named as $L_r$ support vectors since they are selected by the $L_r$ proximal operator. Before proceeding, we first review  the definitions of support vectors of hard margin SVM, hinge loss SVM and ramp loss SVM, respectively.

 {\bf i) Support vectors of hard margin SVM\cite{CV95}:} Let (${\bf\widetilde{w}}; \widetilde{b}$) be a global minimizer of hard margin SVM and  $\bm{\widetilde{\alpha}}=(\widetilde{\alpha}_1,\widetilde{\alpha}_2,\cdots,\widetilde{\alpha}_m)^\top$ with $\widetilde{\alpha}_i\geq0$  be a solution of its dual problem. Then the  global minimizer ${\bf\widetilde{w}}$ satisfies
 \begin{eqnarray}\nonumber
\label{ww1}{\bf\widetilde{w}} =\underset{i \in \widetilde{J}^{*}}{\sum}\widetilde{\alpha}_i y_i\bfx_{i},
\end{eqnarray}
where $\widetilde{J}^{*}:= \{i\in \N_m:\widetilde{\alpha}_i>0\}$. The training vectors $\{\bfx_i, i\in \widetilde{J}^*\}$ are called support vectors. For any $i\in \widetilde{J}^{*}$, we have
 \begin{eqnarray}\nonumber
 y_i(\langle{\bf\widetilde{w}},\bfx_{i}\rangle+\widetilde{b})=1.
\end{eqnarray}

 {\bf ii) Support vectors of hinge loss SVM\cite{CV95}:} Let (${\bf\widehat{w}}; \widehat{b}$) be a  global minimizer of hinge loss SVM and  $\bm{\widehat{\alpha}}=(\widehat{\alpha}_1,\widehat{\alpha}_2,\cdots,\widehat{\alpha}_m)^\top$ with $\widehat{\alpha}_i\in[0, C]$ be a solution of its dual problem. Then the  global minimizer $\bf\widehat{w}$ satisfies
 \begin{eqnarray}\nonumber
\label{ww1}{\bf\widehat{w}} =\underset{i \in \widehat{J}^{*}}{\sum} \widehat{\alpha}_iy_i\bfx_{i},
\end{eqnarray}
where $\widehat{J}^{*}:=\{i\in \N_m:\widehat{\alpha}_i\in(0,C]\}$. The training vectors $\{\bfx_i, i\in \widehat{J}^*\}$ are called support vectors. For any $i\in\widehat{J}^{*}$, we have
 \begin{eqnarray}\nonumber
\begin{cases} y_i(\langle{\bf\widehat{w}},\bfx_{i}\rangle+\widehat{b})=1, ~ i \in \{i\in \N_m:\widehat{\alpha}_i\in(0,C)\},\\
y_i(\langle{\bf\widehat{w}},\bfx_{i}\rangle+\widehat{b})\leq1, ~ i \in \{i\in \N_m:\widehat{\alpha}_i=C\}.
\end{cases}
\end{eqnarray}

{\bf iii) Support vectors of ramp loss SVM\cite{RF06a}:} Let (${\bf\overline{w}}; \overline{b}; \overline{\bfu}$) and $\bm{\overline{\lambda}}\in \R^m$ with $\overline{\lambda}_i\in[-C,0]$ satisfy  KKT conditions. From \eqref{KKT-lam} and the first equation of \eqref{PH-stakkt}, the KKT point ${\bf\overline{w}}$ satisfies
 \begin{eqnarray}\nonumber
\label{ww1}{\bf\overline{w}} =-\underset{i \in \overline{J}^{*}}{\sum} \overline{\lambda}_iy_i\bfx_{i},
\end{eqnarray}
where $\overline{J}^{*}:=\{i\in \N_m:\overline{\lambda}_i\in[-C,0)$. The training vectors $\{\bfx_i, i\in \overline{J}^*\}$ are called support vectors. For any $i\in \overline{J}^{*}$, from \eqref{KKT-lam} and the third equation of \eqref{PH-stakkt}, we have
 \begin{eqnarray}\nonumber
\begin{cases} y_i(\langle{\bf\overline{w}},\bfx_{i}\rangle+\overline{b})\in\{0,1\}, ~ i \in \{i\in \N_m:\overline{\lambda}_i\in(-C,0)\},\\
y_i(\langle{\bf\overline{w}},\bfx_{i}\rangle+\overline{b})\in[0,1], ~ i \in \{i\in \N_m:\overline{\lambda}_i=-C\}.
\end{cases}
\end{eqnarray}

The above results show that the hard margin SVM and hinge loss SVM  define the support vectors at their global minimizer. However, authors in \cite{RF06a} define the support vectors at the KKT point of $L_r$-SVM. In the following, we define the $L_r$ support vectors  for $L_r$-SVM at the P-stationary point, which is also the local minimizer of $L_r$-SVM.

\vskip 2mm
Summarizing the above analysis, we obtain the following interesting result.

\begin{theorem}{\rm\label{support-vector}  Let ($\bfw^{*};b^{*};\bfu^{*}$) with $\bm{\lambda}^*\in \R^m$ and $\gamma>0$ be a P-stationary  point of (\ref{RR-SVM}) for $\gamma C\geq2$. Then all $L_r$ support vectors must fall into the support hyperplanes $\langle \bfw^*,\bfx\rangle+b^*= \pm1$, which possesses the same feature as the one of hard margin SVM.}
\end{theorem}
\section{Conclusions}
%
In this paper, with the help of explicit expression of proximal operator for ramp loss, we have introduced and characterized a novel first-order necessary and sufficient optimality conditions of $L_r$-SVM. We have defined the $L_r$ support vectors based on the concept of P-stationary point and showed that all $L_r$ support vectors fall into the support hyperplanes for $\gamma C\geq2$. Based on the advance of P-stationarity, could we design an efficient algorithm for $L_r$-SVM? We leave this topic to be investigated in the future.

\vskip 2mm
\noindent{\bf\large Acknowledgements}\\

This work is supported by the National Natural Science Foundation of China (11926348-9, 61866010, 11871183), and the Natural Science Foundation of Hainan Province (118QN181).

\renewcommand{\refname}{{\Large References}}

\end{document}